Acrylamide is a chemical that is sometimes found in cooked starchy foods and which is thought to increase the risk of certain kinds of cancer. The paper A Statistical Regression Model for the Estimation of Acrylamide Concentrations in French Fries for Excess Lifetime Cancer Risk Assessmentt describes a study to investigate the effect of frying time (in seconds) and acrylamide concentration (in micrograms per kilogram) in french fries. The data in the accompanying table are approximate values read from a graph that appeared in the paper. FryingAcrylamide 120 270 190 135 (a) If the goal is to learn how acrylamide concentration is related to frying time, which of these two variables is the response variable and which is the predictor variable? O The response variable (y) is the acrylamide concentration, and the predictor variable (x) is the frying time. O The response variable (y) is the frying time, and the predictor variable (x) is the acrylamide concentration. )Construct a scatterplot of these data. (Use the result of (a) for the choice of axis.) 350 300 250 200 150 100 100 150 200 250 300 350 Frying Time (seconds) 100 150 200 250 30 350 Frying Time (seconds) 100 15 200 250 300 350 100 150 200 250 300 350 micrograms per klogram) micrograms per klogram
media%2F9d2%2F9d2ee429-6ed0-42a6-8d40-fd
media%2Fcba%2Fcbadf6a5-3438-4fec-bc0b-4d
Answer #1

a) The response variable Y is the acrylamide concentration and the predictor variable(X) is the frying time.

B) the scatter plot in the upper right corner is the correct one.

Here horizontally frying time has been there as X axis.

C) I have done this in R.

R code-

X=c(150,240,240,270,300,300)
Y=c(155,120,195,190,135,270)
lm(Y~X)

Output-

Coefficients:
(Intercept)            X  
    90.9615       0.3462

So, Y_hat= 90.962 + 0.346*X

D) The slope=0.346 is positive and it is consistent with the observed association...

As per HOMEWORKLIB RULES we should solve only the first four sub parts of the question. For the other parts please post the question differently mentioning your requirement. Thank you